OpenStudy (anonymous):

In the data set, n is an outlier and is the greatest number in the data set. 10 19 21 24 25 29 30 34 37 n What is the least possible whole number value for n? @ganeshie8

ganeshie8 (ganeshie8):

let \(n\) be some large number and find the `median` and `IQR`

OpenStudy (anonymous):

@ganeshie8 How do I find the median in an even amount of numbers, I can't remember?

ganeshie8 (ganeshie8):

take the average of middle two values

ganeshie8 (ganeshie8):

10 19 21 24 `25 29` 30 34 37 n median (25+29)/2 = ?

OpenStudy (anonymous):

Oh, uh... 27

ganeshie8 (ganeshie8):

Right! next find the IQR

ganeshie8 (ganeshie8):

IQR = Q3 - Q1

ganeshie8 (ganeshie8):

do you know how to find Q1, Q3 ?

OpenStudy (anonymous):

21 and 30?

ganeshie8 (ganeshie8):

Q1 = 21 is correct Q3 is wrong

ganeshie8 (ganeshie8):

`10 19 21 24 25` `29 30 34 37 n` clearly 21 is the center of first half, so Q1 = 21 similarly 34 is the center of second half, so Q3 = 34

ganeshie8 (ganeshie8):

next find the IQR

ganeshie8 (ganeshie8):

sim[ply subtract Q1 from Q3 for IQR : \[IQR = Q3-Q1 = 34-21 = 13\]

ganeshie8 (ganeshie8):

we call a data value is outlier if it is greater than `median + 1.5* IQR ` median = 27 IQR = 13 `median + 1.5*IQR` = 27 + 1.5*13 = 46.5 so the minimum possible whole nuumber value for n is 47 for it to be an outlier

OpenStudy (anonymous):

So the answer is 47?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

@ganeshie8 47 turned out to be wrong, it was 54. Thanks though! :)